### v2

# Bridging Compositional and Distributional Semantics: A Survey on Latent Semantic Geometry via AutoEncoder

Abstract
Integrating compositional and symbolic properties into current distributional semantic spaces can enhance the interpretability, controllability, compositionality, and generalisation capabilities of Transformer-based auto-regressive language models (LMs). In this survey, we offer a novel perspective on latent space geometry through the lens of compositional semantics, a direction we refer to as semantic representation learning. This direction enables a bridge between symbolic and distributional semantics, helping to mitigate the gap between them. We review and compare three mainstream autoencoder architectures— Variational AutoEncoder (VAE), Vector Quantised VAE (VQVAE), and Sparse AutoEncoder (SAE)—and examine the distinctive latent geometries they induce in relation to semantic structure and interpretability.
1 Introduction
Representation Learning.
Representation learning encompasses a broad set of techniques aimed at automatically discovering useful features and structures from complex, high-dimensional data. The core objective is to transform raw inputs into compact, informative representations that capture the underlying geometry, patterns, and semantic relationships within the data Bengio et al. (2013); Alshammari et al. (2025). By learning to encode the essential factors of variation, representation learning enables models to generalise more effectively and supports interpretability and modularity across a wide range of domains, including vision Radford et al. (2021), speech Liao et al. (2024), language Yang et al. (2025a), and multimodal learning Yang et al. (2025b); Liu et al. (2025).
Compositional Semantics.
In the field of natural language learning and generation, improving the geometric structure of semantic features within the representation space can enhance the controllability and interpretability of language models (LMs), which has been widely explored in the task, such as style-transfer Liu et al. (2023); Gu et al. (2023, 2022); Hu and Li (2021); Vasilakes et al. (2022). Those studies provide a quasi-symbolic, localised, compositional generation control to informal, less-rigid semantic features, such as sentiment and topic. Recently, a line of studies investigated this kind of control over formal semantic features Zhang et al. (2024b, 2025a, e, c, d, 2025b). Formal/symbolic semantics, such as Montague Semantics Dowty et al. (2012), Semantic Role Labelling Palmer et al. (2010), and Argument Structure Theory Jackendoff (1992), which provides a canonical, granular, and rigid representation of meaning in natural language, has been a central focus of philosophical and linguistic inquiry for thousands of years. Its goal is to systematically capture the logical structure of linguistic expressions and to model how complex meanings arise from simpler parts. All these studies provide valuable insights to bridge the compositional and distributional semantics.
Language Models.
Over the last few years, Transformer-based auto-regressive LMs with distributional semantic space have emerged as powerful tools that offer a flexible and scalable foundation for addressing a broad and diverse spectrum of natural language tasks Touvron et al. (2023). However, despite their empirical success, they often operate as opaque black boxes with limited insight into the underlying semantic mechanisms they use. This raises an important question: can we develop language representations or models that incorporate the properties of symbolic, structural, and compositional semantics? Addressing this question yields several important advantages:
Interpretability and Controllability: Models grounded in symbolic and structured representations are inherently more controllable. For example, in the sentence: animals require oxygen for survival, the words are functionally combined into sentence semantics: $\lambda x(\text{animals}(x)\rightarrow\text{require}(x,\text{oxygen}))$ where $x$ is the variable of any entity within a logical structure. In this case, we can localise the sentence semantics by replacing $x$ with birds, etc. This localised process has been applied in Natural Logic, which manipulates the surface form of sentences during inference Bowman et al. (2015); Angeli and Manning (2014); MacCartney and Manning (2007).
Safety and Trustworthiness: As language models are increasingly deployed in high-stakes domains, ensuring their safety and trustworthiness becomes critical. Systems that incorporate formal semantic structures are more likely to produce outputs that are logically coherent and consistent with intended meanings. Symbolic reasoning frameworks can help verify that a system adheres to specified rules or constraints, reducing the likelihood of hallucinations, contradictions, or harmful outputs. This alignment with formal systems of logic and meaning contributes to the reliability and accountability of AI systems Chen et al. (2024).
Composition and Generalisation: One of the enduring strengths of compositional semantics is its ability to handle compositionality—the principle that the meaning of a complex expression is determined by the meanings of its parts and the rules used to combine them. Embedding this principle into language models enhances their ability to generalise to novel combinations and linguistic constructions, beyond those seen during training. In the Math Reasoning task Meadows et al. (2024), for example, the model is tasked with predicting the output of a mathematical transformation given a specified operation and an input formula during training. However, current LMs often fail to generalise effectively to variable or length out-of-distribution (OOD) inputs Zhang et al. (2024e). Structured semantic representations—such as syntax trees that explicitly encode the function, constant, and variable—can provide a more robust framework for generalisation, enabling models to maintain compositional integrity even in the face of lexical variation.
Our Scope.
To integrate these potential advantages into current LM, this survey provides a systemic review of studies that focus on explicitly embedding symbolic, compositional, and semantic properties into the latent space of LMs - a direction we term semantic representation learning. We seek to bridge the gap between symbolic semantics and neural representations, facilitating more interpretable and controllable to LMs’ generation. The survey is organised as follows: (a) Section 2 frames the scope of the survey, stating a definition of semantic representation learning; (b) Section 3 reviews the main data resources for tasks requiring compositional semantics; (c) Section 4 provides a detailed classification of the main architectural patterns and approaches proposed for semantic representation learning; (d) Section 5 describes quantitative and qualitative metrics for the evaluation of semantic representation space; (e) Section 6 outlines open challenges and promising directions for future research.
Expression Compositional Semantic Tasks Datasets & Description
Definitions Definition Modelling(vec2def) Noraset et al. (2016), Yang et al. (2019), Mickus et al. (2022)
Definition Modelling(word2def) Noraset et al. (2016), WordNet, Wiktionary, Wikipedia
Reverse Dictionary(def2vec) Semeval-2022 Task 1: CODWOE Mickus et al. (2022)
Explanations QA with explanation graph WorldTree Jansen et al. (2018)
QA with entailment tree EntailmentBank Dalvi et al. (2021)
Syllogistic NLI Rule-based Explanatory NLI corpus Zhang et al. (2024d)
commonsense QA Zebra-KB Molfese et al. (2024)
explanations corpora survey Thayaparan et al. (2020)
Others Sentiment Style-transfer Yelp reviews Shen et al. (2017), Yahoo answers Yang et al. (2017)
IMDb movie review Diao et al. (2014); Maas et al. (2011); Li et al. (2018)
Stanford Sentiment Treebank Socher et al. (2013)
Formality Style-transfer GYAFC Rao and Tetreault (2018)
Codeswitching Style-transfer LinCE Aguilar et al. (2020)
Topic Style-transfer AGNews dataset Zhang et al. (2015), DBpedia Zhang et al. (2015)
Toxic Style-transfer Jigsaw Toxic Comment
Disentanglement dSentences
style-transfer corpora survey Jin et al. (2022)
Table 1: Summarisation of data sources and tasks. Please note that our motivation is to investigate the utilisation of those expressions in fundamental representation learning, rather than the downstream tasks we mentioned here.
2 Semantic Representation Learning
In this work, we define semantic representation learning as the process of encoding compositional, symbolic semantic properties into the latent space of language models in a way that supports both interpretability and controllability. Specifically, we characterise this through two key dimensions:
The Scope of Interpretability.
One promising direction to interpret the representation space geometry is Linear Representation Hypothesis, which suggests that high-level concepts can be encoded within causally separated subspaces (i.e., directions) of the high-dimensional latent space. This hypothesis has been supported across various domains, including word embeddings (Mikolov et al., 2013; Arora et al., 2016), sentence embeddings (Li et al., 2020a), and Transformer-based language models (Meng et al., 2022; Merullo et al., 2024; Chanin et al., 2024; Park et al., 2024a, b).
In contrast to prior work, we emphasise that a well-structured representation space should exhibit interpretable geometric properties—such as meaningful directions, distances, and clustering—that correspond to underlying “semantic” features. For instance, semantically similar expressions should be mapped to nearby vectors, while distinct semantic roles or categories should occupy separable regions or align along distinct axes in the latent space. Such geometric interpretability offers valuable insights into how language models internally organise and encode linguistic meaning.
The Scope of Controllability.
By manipulating these interpretable geometric structures, it should be possible to exert fine-grained, quasi-symbolic, localised generation control over the model. This includes altering specific semantic attributes (e.g., tense, sentiment, or argument structure) by adjusting latent representations, thus enabling more precise and targeted language generation. For example, modifying sentiment from I love this movie to I hate this movie, or adjusting the argument of semantic role from animal is a kind of living thing to tiger is a kind of living thing.
3 Data Resources
In this section, we primarily introduce two types of natural language expressions that are widely employed in tasks involving compositional formal semantics. Both types of data exhibit semantically rich yet structurally controlled properties, making them suitable for evaluating the compositional capabilities of representation spaces. In addition, we briefly discuss other resources commonly used to study informal semantics [^1] In this work, we distinguish the informal and formal semantics by considering whether the latent space can explicitly encode the structural representation, such as syntax, semantic roles, etc..
Natural Language Definitions.
Natural language definitions are linguistic expressions that explain the meaning of terms using everyday language, allowing individuals to comprehend, categorise, and communicate abstract concepts. They serve as bridges between raw lexical forms and the conceptual knowledge they denote, facilitating understanding across contexts and domains. These definitions are often characterised by a high level of abstraction and structured phrasing, which mirrors the hierarchical organisation of concepts in human cognition. This structured nature plays a key role in supporting reasoning, knowledge transfer, and the construction of mental models, making definitions an essential component of both human learning and machine understanding.
Natural language definitions are widely applied in the task, such as definition modelling Noraset et al. (2016) where the goal is to generate a natural language definition given the word to be defined (definiendum) or its word embedding, and reversed dictionary (definition-to-vector) Mickus et al. (2022). These tasks are closely connected to compositional semantics, as they require the model to parse, interpret, and reconstruct meaning based on the combination of words and structures in the input or output. By grounding lexical meaning in natural language, these approaches contribute to more interpretable and semantically aware representation learning.
Natural Language Explanations.
Natural language explanations, such as atomic sentence: animal is a kind of living thing and condition sentence: if something is a living thing, it needs water to survival, are designed for formal/compositional/symbolic semantic inference task in natural language form, which provides a semantically complex and yet controlled experimental setting, containing a both well-scoped and diverse set of target concepts and sentence structures. A number of scientific or commonsense reasoning benchmarks leverage natural language explanations as core resources, including WorldTree Jansen et al. (2018), EntailmentBank Dalvi et al. (2021), and Zebra Molfese et al. (2024).
Other Resources.
In addition to the two main types of expressions we focus on in compositional formal semantics, the NLP domain includes a wide range of corpora that explore compositional informal semantics across various downstream tasks, such as style transfer and paraphrasing. These datasets often capture nuanced variations in meaning associated with attributes such as sentiment, formality, or tone, and serve as valuable resources for investigating how language models handle semantic composition in more flexible and context-dependent settings. Typical corpora include Yelp reviews Shen et al. (2017), Yahoo answers Yang et al. (2017), dSentences  [^2] https://github.com/mcharrak/discreteVAE, IMDb movie review Maas et al. (2011), AGNews dataset Zhang et al. (2015), and Jigsaw Toxic Comment Classification Challenge Dataset  [^3] https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge/. More style-transfer corpora can be found in the survey Jin et al. (2022).
{forest}
for tree= grow=east, draw, rounded corners, node options=align=center, font=, edge=->, >=latex, parent anchor=east, child anchor=west, l sep+=10pt, s sep+=5pt, anchor=base west, [Latent Semantic Geometry [Sparse AutoEncoder, fill=yellow!50 [Informal Linguistic Attribute Space, fill=yellow!20 [Elhage et al. (2022); Cunningham et al. (2023); Deng et al. (2025); He et al. (2025)] ] [Formal Linguistic Attribute Space, fill=yellow!20 [Jing et al. (2025)] ] ] [Vector Quantised VAE, fill=cyan!60 [Semantic or Concept Space, fill=cyan!30 [Zhang et al. (2024c); Garg et al. (2025)*] ] [Syntax-Semantic Space, fill=cyan!30 [Hosking and Lapata (2021)] ] [Hierarchical Syntactic Space, fill=cyan!30 [Hosking et al. (2022)] ] [Topic Attribute Space, fill=cyan!30 [Yoo and Choi (2024)] ] ] [Variational AutoEncoder (VAE), fill=green!40 [Reasoning Attribute Space, fill=green!20 [Natural Language, fill=green!10 [Zhang et al. (2024d, 2025b)] ] [Others (e.g., Vision), fill=green!10 [Sun et al. (2018); Bonnet and Macfarlane (2024); van Krieken et al. (2025a)] ] ] [Semantic Attribute Space, fill=green!20 [Carvalho et al. (2023); Zhang et al. (2024b); Zhang et al. (2025a)] ] [Syntax-Semantic Attribute Space, fill=green!20 [Zhang et al. (2019); Bao et al. (2019); Huang et al. (2021); Chen et al. (2019); Zhang et al. (2024e)] ] [Syntax Attribute Space, fill=green!20 [Shi et al. (2021); Huang and Chang (2021); Mercatali and Freitas (2021); Felhi et al. (2022)] ] [Style-transfer Attribute Space, fill=green!20 [Single Attribute, fill=green!10 [Fang et al. (2019); John et al. (2019); Hu et al. (2017); Xu et al. (2021); Mai et al. (2020);
Huang et al. (2020); Li et al. (2022); Yoshioka et al. (2022); Liu et al. (2020, 2023);
Duan et al. (2020); Shen et al. (2020); Zhang et al. (2024a)*] ] [Multiple Attributes, fill=green!10 [Gu et al. (2023, 2022); Hu and Li (2021); Vasilakes et al. (2022)] ] ] ] ]
Table 2: Taxonomy for latent semantic geometry via autoencoder architecture. For VAE and VQVAE, * indicates configurations that are plug-and-play inside LMs, rather than considering the LM as a single decoder.
4 Approach
This section outlines the major architectural trends for learning semantic representation spaces and their integration with language models.
4.1 Sentence-level Semantic Space
The Variational Autoencoder (VAE) Kingma and Welling (2013) is a generative model that learns a continuous and probabilistic latent space through an encoder-decoder architecture. The encoder maps input data to a distribution over latent variables—typically modeled as a multivariate Gaussian—while the decoder reconstructs the original input from samples drawn from this latent distribution. By jointly optimising a reconstruction loss and a regularization term (the Kullback-Leibler divergence between the approximate posterior and a prior distribution), the VAE encourages the latent space to be smooth and structured, making it well-suited for capturing semantic variations in a principled and interpretable way.
4.1.1 Transformer Language VAEs
With the advancement of Transformer-based LMs, recent Language VAEs commonly adopt Transformer-based pretrained models as both encoder and decoder components. Notable examples include Optimus Li et al. (2020b), Della Hu et al. (2022), AdaVAE Tu et al. (2022), LlamaVAE Zhang et al. (2023), and a general Language VAE framework designed to support flexible variable combinations between encoder and decoder Carvalho et al. (2025).
Although more complex latent space manifolds—such as those employed in hyperbolic VAEs Cho et al. (2023) and hyperspherical VAEs Davidson et al. (2018)—have been proposed to encode different geometrical properties, the majority of Transformer-based language VAE models continue to adopt a Gaussian distribution as the prior over the latent space, due to its mathematical tractability, ease of optimisation, and its suitability for modeling continuous, smooth variations in semantics. The Gaussian prior also facilitates effective sampling and interpolation within the latent space, which is essential for tasks like conditional generation and semantic manipulation.
Latent Space Injection.
One representative VAE-based model is Optimus Li et al. (2020b), which employs BERT Devlin et al. (2018) as the encoder and GPT-2 Radford et al. (2019) as the decoder. The model is designed to learn rich latent representations of sentences by integrating pretrained language models into the VAE framework.
Specifically, BERT encodes the input sentence $x$ and produces a fixed-length representation using the special [CLS] token. This representation is used to parameterize a Gaussian latent distribution $\mathcal{N}(\mu,\Sigma)$, where both $\mu$ and $\Sigma$ are learned through training. A latent vector $z\sim\mathcal{N}(\mu,\Sigma)$ is then sampled and passed through a multi-layer perceptron $W$, which projects $z$ into a high-dimensional embedding $h\in\mathbb{R}^{D\times L\times H}$. Here, $D$, $L$, and $H$ denote the number of attention heads, the sequence length, and the hidden size, respectively.
(1) KV Memory Injection. The projected $z$ are then incorporated as additional keys and values in the self-attention mechanism of each Transformer layer in the decoder—a configuration referred to as the memory setup. This design allows the latent variable to influence every layer of the GPT-2 decoder, enabling fine-grained control over generation. The attention operation at each layer can be summarised as:
[…]
where $Q$, $K$, and $V$ represent the query, key, and value matrices commonly used in the Transformer architecture Vaswani et al. (2017).
(2) KV Addition Injection. In addition to memory setup, $z$ can also be directly incorporated into the key and value matrices of the self-attention mechanism, as demonstrated in Zhang et al. (2024e).
[…]
(3) Query Addition Injection. $z$ can also be directly added into the Query matrices of the self-attention mechanism for effectively search the pretrained, memorised knowledge from Key and Value by incorporating prior information Morris et al. (2025). For example, Zhang et al. (2024e) inject syntax information into Query for effective searching the semantic information from Key and Value.
[…]
(4) KV Fusion Injection. Moreover, Della Hu et al. (2022) adapted the tensor fusion operation Liu et al. (2018). In more detail, given a hyper-parameter, rank $r=4$, the $\tilde{K}$ can be described as:
$\tilde{K}=(\sum_{i=1}^{r}W_{k}^{i}[K;\mathbbm{1}])\circ(\sum_{i=1}^{r}W_{z}^{i}[z;\mathbbm{1}])$ (4)
Where $\mathbbm{1}$ is the matrix of ones, $W_{z}^{i}$ and $W_{k}$ are the trainable linear transformations. Same operation for Value. The final softmax function can be described as:
[…]
(5) KV Cache Injection. In the previous setups, training a VAE model requires fine-tuning the decoder in an end-to-end manner. However, the emergence of large language models (LLMs) has made such pre-training significantly more challenging due to their scale and complexity. To address this issue, recently, Carvalho et al. (2025) propose a novel injection mechanism that enables more efficient integration of latent variables without the need for full decoder fine-tuning.
They use the KV caching mechanism of the Causal LM model classes within the transformers library to inject a positional projection of the latent vector. A linear projection of the latent vector $h_{cache}=W_{m}z$ plays the role of an additional context to guide generation, in the form of hidden KV cache entries $X^{h}_{t}$ interleaved with those produced by the decoder, where $W_{m}\in\mathbb{R}^{LH\times S}$ is separated into $S\times L$ (sequence length * # layers) vectors of hidden size $H=K\times V$.
Latent injection setup & Math formula Summarisation
KV Memory: softmax(Q [z;K]Td) [z;V] LM decoder only rely on attached token $z$ which is more likely to be ignored by self-attention, even exacerbating KL vanishing Li et al. (2020b).
KV Addition: softmax(Q [z + K]Td) [z + V] Simple and global interventions in the sequence of attention of LM decoder, which can mitigate the KL vanishing without additional parameters Zhang et al. (2024e).
Query Addition: softmax([Q+z] KTd) V Integrating the latent representation $z$ into decoder via query can deliver effective searching over Key and Value in decoders’ parametric space Zhong et al. (2025).
KV Fusion: ^K = (∑_i=1^r W^i_k [K;1]) ∘(∑_i=1^r W^i_z [z;1]) Using tensor fuse operation to deeply fuse latent variables with hidden states in the LM decoder, but requiring additional training parameters Hu et al. (2022).
KV Cache: h_cache = W_m z Flexibly integrating $z$ into the attention of various LM decoders with the cost of perturbing pretrained knowledge and more training parameters Carvalho et al. (2025).
Latent Prompt: P = MLP(z) (where $P$ are $k$ soft prompt embeddings) Seamlessly and flexibly integrating $z$ into LM decoders, but work on high-dimensional token embeddings. High dimensionality and complexity of the sequence space makes the search/optimisation difficult Zhou et al. (2024).
Table 3: Latent space injection methods in Transformer-based Language VAE.
(6) Latent Prompt Injection. However, KV cache injection setup requires the replacement of attention representations, which can interfere with the internal knowledge encoded in the LLM. To seamlessly integrate latent variables while preserving the LLM’s pretrained knowledge, alternative strategies are needed. Zhou et al. (2024) consider the VAE latent representation as the prompt of LLMs. Specifically, after obtaining the latent representation $z$, they use an upper MLP to map it into k soft prompt token embeddings. These soft embeddings serve as a steering vector, which is concatenated before the <BOS> token to assist the LLM in decoding.
Figure 1: VAE architecture where the latent space $z$ can present different geometrical properties.
Figure 2: VQVAE architecture where the latent space $z$ can present different geometrical properties.
Figure 3: SAE architecture where the latent space $z$ can present different geometrical properties.
4.1.2 Latent Space Geometry
According to their distinct geometrical properties, we categorise these approaches into the following types of semantic representation spaces:
(1) Style-transfer Attribute Space. Style transfer tasks aim to modify sentences with respect to specific informal linguistic attributes—such as sentiment, formality, or affirmation/negation—while preserving the core semantic content. A number of studies have focused on sentiment control, learning latent space to isolate and manipulate emotional tone John et al. (2019); Li et al. (2022); Liu et al. (2020). Other works target multi-attribute control, where models are trained to simultaneously manage multiple stylistic dimensions within the latent space Liu et al. (2023); Gu et al. (2023, 2022); Hu and Li (2021); Vasilakes et al. (2022). These approaches leverage separated feature representations to enable fine-grained and interpretable text transformations.
Next, we introduce four types of latent space geometries that are designed to encode more formal semantic features (such as syntax, semantic role label, and reasoning logics).
(2) Syntax Attribute Space. Syntactic roles can be considered as generative factors within the latent space of VAEs, where each factor captures a distinct aspect of sentence structure or grammatical function Mercatali and Freitas (2021); Felhi et al. (2022). By disentangling these roles in the latent representation, such models aim to isolate syntactic information (e.g., subject, object, modifier) from lexical or semantic content, thereby enabling more structured and interpretable generation and manipulation of language.
(3) Syntax-Semantic Attribute Space. Several studies have explored methods for learning latent spaces in which syntactic structure and lexical semantics are disentangled. For example, Zhang et al. (2019) used two separate encoders to independently encode semantic and syntactic information. In contrast, Bao et al. (2019) utilised a single encoder within a VAE framework to disentangle semantic and syntactic representations through multi-task optimisation. Additionally, Zhang et al. (2024e) employed a graph-based encoder alongside a semantic encoder to explicitly capture syntax and semantics, mitigating the information bottleneck from a single encoder.
(4) Semantic Attribute Space. Defining generative factors in terms of semantic roles has recently been explored as a promising direction for structuring latent spaces. Carvalho et al. (2023) defined the generative factors of natural language definition, named Definition Semantic Roles (DSR) Silva et al. (2018), to assist the Definition Modelling tasks. Recent studies Zhang et al. (2024b, 2025a) formalise the latent space geometry from the perspective of Argument Structure Theory Jackendoff (1992); Levin (1993); Rappaport Hovav and Levin (2008). In this framework, each semantic role–word content combination is represented as a convex cone within the latent space, enabling localised semantic control and enhancing interpretability.
(5) Reasoning Attribute Space. More recent studies have also explored the encoding of reasoning patterns within the latent space of VAEs for Natural Language Inference (NLI) tasks Zhang et al. (2024d, 2025b). They seek to structure the latent space in a way that captures inferential relationships, such as mathematical reasoning rules, thereby enabling the model to represent not only the meanings of individual sentences but also the logical relations that hold between them.
4.2 Token-level Semantic Space
4.2.1 Vector Quantized VAEs
Overview.
The Vector Quantized Variational Autoencoder (VQ-VAE) Van Den Oord et al. (2017) is a variant of the traditional VAE that replaces the continuous latent space with a discrete latent codebook. Rather than sampling from a Gaussian prior, the encoder outputs are mapped to the nearest vector in a learnable set of embedding vectors, known as the codebook. This discrete codebook serves as a fundamental representation space that can be efficiently transferred across architectures. As a result, VQ-VAE has been widely adopted in domains such as speech (e.g., Fish-Speech Liao et al. (2024)) and multimodal learning (e.g., Janus-Pro Chen et al. (2025)).
Architecturally, the model comprises three main components: an encoder that transforms input data $x$ into continuous latent vectors, a quantization module that replaces each vector with its nearest neighbor in the codebook, and a decoder that reconstructs the input from these quantized representations. Training involves minimizing a combination of reconstruction loss, a vector quantization loss to align encoder outputs with codebook entries, and a commitment loss that encourages encoder outputs to stay close to selected codebook vectors. Unlike VAEs, there is no standardized architecture for language VQ-VAEs. Most existing models are custom-designed to specific downstream tasks. As a result, we do not provide a comprehensive overview of these individual implementations in this survey.
Latent Space Geometry.
Recent extensions of the VQ-VAE framework have demonstrated its effectiveness in learning discrete latent semantic spaces tailored to specific NLP tasks, delivering different geometrical properties.
(1) Topic Attribute Space. Topic-VQVAE Yoo and Choi (2024) focuses on document-level generation by learning topic-aware latent codes. It leverages vector quantization to discover discrete topic representations that guide the generation process, enabling more coherent and thematically consistent document outputs.
(2) Hierarchical Syntactic Space. Natural language exhibits a hierarchical structure across multiple linguistic levels, particularly at the syntactic level, where constituents recursively contain sub-constituents. To capture this property, HRQ-VAE Hosking et al. (2022) introduces a hierarchical syntactic sketch space, which represents syntax in a nested, discrete latent structure. This approach enables the model to better reflect the recursive and compositional nature of syntactic patterns in language generation and understanding.
(3) Syntax-Semantic Space. Hosking and Lapata (2021) propose a model that factorizes sentence meaning into discrete syntactic and continuous semantic components. Focusing on the task of question paraphrasing, their model reconstructs a question by combining the surface form from an exemplar with the semantic content from a paraphrase. This dual-space approach enables more controlled generation, allowing for explicit manipulation of syntax and semantics independently while preserving overall meaning.
(4) Semantic Space. To investigate the geometrical properties of discrete latent spaces, a recent study by Zhang et al. (2024c) integrates the VQ-VAE framework into the T5 encoder-decoder architecture to learn a discrete semantic space. The authors further propose techniques such as traversal and interpolation to explore and manipulate representations within this discrete space, targetting fine-grained semantic control.
4.2.2 Sparse AutoEncoder
Overview.
Previous studies have revealed that LMs often exhibit polysemanticity, where a single neuron responds to multiple, semantically unrelated inputs Elhage et al. (2022). This entanglement of concepts within individual neurons poses a significant challenge to interpretability and fine-grained control. Sparse Autoencoders (SAEs) have emerged as a promising approach to mitigate this issue. SAEs are designed to learn a sparse, linear, and decomposable representation of the internal activations of LMs, where each dimension—termed a base vector—ideally corresponds to a distinct, interpretable linguistic feature. By enforcing a sparsity constraint—ensuring that only a small number of features are active at any given time—the model encourages each active feature to align with a distinct and interpretable concept Yun et al. (2021); Cunningham et al. (2023).
Architecturally, a sparse autoencoder consists of a linear encoder that maps input representation to a higher dimensional feature space and a linear decoder that reconstructs the input from this representation. To promote sparsity, an additional regularization term—such as Kullback-Leibler (KL) divergence or an L1 penalty—is applied to the hidden activations, encouraging most neurons to remain inactive (i.e., output values close to zero) for any given input. This constraint forces the model to learn meaningful, distinct features and prevents trivial identity mappings. A comprehensive survey of SAEs, including architectural variants and optimisation strategies, is provided by Shu et al. (2025).
Latent Space Geometry.
Recent studies have shown that SAEs exhibit meaningful linguistic geometrical properties, which can be leveraged to control language model generation. We categories the geometrical properties into two main types.
(1) Informal Linguistic Attribute Space. This refers to the organisation of latent dimensions corresponding to loosely defined or stylistic language properties. SAEs have been shown to uncover interpretable features aligned with lexical semantic attributes, enabling controlled generation and feature-level interventions Elhage et al. (2022); Cunningham et al. (2023). Additionally, recent work has investigated the behavior of LLMs in multilingual settings, revealing that SAEs can also capture language-specific features and align them across languages in the latent space Deng et al. (2025). Recent work further investigates task-specific attribute separation in open-ended generation control, such as sentiment, truthfulness, and politics polarity steering tasks He et al. (2025).
(2) Formal Linguistic Attribute Space. Recent work investigated the linguistic capabilities of LLMs Jing et al. (2025). The authors extract fine-grained linguistic features across six dimensions—phonetics, phonology, morphology, syntax, semantics, and pragmatics—by encoding LLM hidden states into sparse, interpretable representations. They evaluate these features using custom datasets, including minimal contrast and counterfactual sentence pairs, and define two metrics: Feature Representation Confidence (FRC) and Feature Intervention Confidence (FIC) to assess how well features correspond to and control linguistic phenomena. Through targeted interventions, they demonstrate that specific features causally affect LLM outputs, providing insights into deep semantic processing and cross-layer linguistic representation.
5 Geometrical Evaluation
Category Metrics Summarisation
Dimension-based $z_{min\_var}~{}error$ Kim and Mnih (2018) Identify dimensions with the smallest variance across different samples for a given factor.
Modularity Ridgeway and Mozer (2018) Assess how exclusively a dimension is associated with a single factor. If MI for one factor dominates a dimension and others are close to zero, modularity is high.
Disentanglement Score Eastwood and Williams (2018) Use feature importances from a classifier to compute entropy for each dimension; lower entropy = better disentanglement.
Factor-based Mutual Information Gap Chen et al. (2018) Measure how much more one latent dimension knows about a factor compared to others. A larger gap between the top two mutual information (MI) scores for a factor indicates stronger disentanglement.
Completeness Score Eastwood and Williams (2018) Similar to Disentanglement Score, but entropy is computed over factors rather than dimensions.
Informativeness Score Eastwood and Williams (2018) Use a classifier to predict factor labels from latent codes; score is based on prediction error (lower error = higher informativeness).
Explicitness Ridgeway and Mozer (2018) Determine how linearly separable the factor information is from the latent representation. Use ROC AUC to quantify separability.
Table 4: Latent space evaluation via disentanglement metrics which quantify the feature-dimension alignment.
Category Approach Summarisation
VAE Steering style-transfer Mercatali and Freitas (2021) given two lists of sentence vectors with contrast generative factors, such as positive and negative, calculating the direction from one to another via subtraction: $\text{vec}_{positive\rightarrow negative}$.
Formal semantic control Zhang et al. (2025a) choosing two sentence vectors within the same semantic role-word content, the added sentence holds the same role-content.
LM Steering CAA Rimsky et al. (2024) Adding the mean activation difference between positive and negative examples during inference to steer model outputs.
RePe Zou et al. (2023) Perturbing activations along principal components of class-conditional differences.
ITI Li et al. (2023) Shifting attention head activations during inference along truth-related directions found via linear probing.
Top PC Im and Li (2025) Projecting activations onto the first principal component of the embedding space, capturing the direction of maximal variance.
SAE Steering SAE-SSV He et al. (2025) Training linear classifiers to identify a small subspace of task-relevant dimensions in latent representations.
SPARE Zhao et al. (2025) Identifying and applying the functional features of knowledge selection behaviours to edit the internal activations of LLMs.
Table 5: Geometrical control and evaluation via latent arithmetic.
The geometry of the latent space can be evaluated from two complementary perspectives: visual inspection and quantitative disentanglement analysis.
5.1 Latent Space Evaluation
Latent Visualisation.
The structure of the latent space can be examined by visualizing feature clusters and their separability. Dimensionality reduction techniques such as Principal Component Analysis (PCA), t-SNE Van der Maaten and Hinton (2008), and UMAP McInnes et al. (2018) are commonly employed to project high-dimensional latent representations into two or three dimensions. These visualisations help reveal semantic groupings, continuity, and separation across different latent factors.
Disentanglement Metrics.
VAE has proven remarkably effective for disentangled representation learning Bhowal et al. (2024). In this case, disentanglement metrics have been proposed to quantify the dimensional separability and feature-dimension alignment within the latent space of VAE. They can be broadly classified into two categories: dimension-based and factor-based metrics.
(1) Dimension-based metrics assess disentanglement by analysing how the values of a single latent dimension change when different generative factors vary. These metrics measure whether a given dimension remains stable across unrelated factors, indicating that it captures a specific, isolated semantic feature. Commonly used metrics include $z_{min\_var}~{}error$ Kim and Mnih (2018), Modularity Ridgeway and Mozer (2018), Disentanglement Score Eastwood and Williams (2018).
(2) Factor-based metrics evaluate disentanglement from the perspective of individual generative factors. They measure how multiple latent dimensions respond when a single factor is varied, assessing whether that factor is encoded compactly and consistently across the latent space. Those metrics include Mutual Information Gap ($MIG$Chen et al. (2018), Completeness Score Eastwood and Williams (2018), Informativeness Score Eastwood and Williams (2018), Explicitness Ridgeway and Mozer (2018).
All metrics statistically measure the feature-dimension alignment. In what follows, we provide a detailed illustration of one representative metric to demonstrate how such alignment is quantified in practice. The $z_{\text{min\_var}}$ error Kim and Mnih (2018) focuses on evaluating how well individual latent dimensions correspond to a specific generative factor. The procedure involves randomly selecting a set of sentences and extracting their latent representations. For each generative factor (e.g., sentiment, tense), the variance of each latent dimension is measured across the selected samples. If one latent dimension exhibits significantly lower variance than the others for that factor, it suggests that this dimension consistently encodes that factor—indicating a higher degree of disentanglement in the latent space. All of the aforementioned metrics have been implemented in a publicly available repository Carvalho et al. (2025) [^4] https://github.com/neuro-symbolic-ai/LangSpace, providing accessible tools for evaluating disentanglement and structure in latent spaces.
5.2 Generation Evaluation
Latent Traversal.
In VAE and VQVAE, latent traversal is a standard procedure for evaluating the geometry of latent spaces (Higgins et al., 2017; Kim and Mnih, 2018). It involves performing a random walk, such as Ornstein-Uhlenbeck, across the latent space. If the latent space exhibits well-structured geometrical properties, the traversed examples are expected to display consistent and uniform patterns.
Latent Interpolation.
The interpolation mechanism operates by selecting two input sentences, $x_{1}$ and $x_{2}$, and encoding them into their corresponding latent representations $z_{1}$ and $z_{2}$. A linear interpolation is then performed along the latent space trajectory defined by $z_{t}=z_{1}\cdot(1-t)+z_{2}\cdot t$, where $t$ varies from 0 to 1 in increments of 0.1. Consequently, nine intermediate sentences are generated along the interpolation path. Prior work has shown that improved latent space geometry leads to smoother and more semantically coherent interpolations Zhang et al. (2024c).
Latent Arithmetic.
Latent arithmetic, such as addition and subtraction, has been widely regarded as a lightweight approach to controlling LM, indicating the geometrical separation and linearity in the latent or parametric space.
In VAE, for example, given two sentences that share the concept animal—animals require food for survival and animals require warmth for survival—their vector addition in the latent space may yield a new sentence such as animals produce milk, which also preserves the concept animal. This suggests the existence of a “convex cone” corresponding to the semantic concept animal within the latent space Zhang et al. (2024c, 2025a). Moreover, using arithmetic to steer the task-specific generative factor, such as sentiment, has been widely deployed in language VAE domain Shen et al. (2020); Mercatali and Freitas (2021). Furthermore, controlling (Large) LMs generation through the arithmetic of steering vectors to internal representations has been widely explored Turner et al. (2023), such as Concept Activation Addition (CAA) Rimsky et al. (2024), Representation Perturbation (RePe) Zou et al. (2023), Top PC Im and Li (2025), and Inference-Time Intervention (ITI) Li et al. (2023). In this context, the discrete concept vectors learned by SAEs can be interpreted as steering vectors, enabling targeted modulation of language model outputs Chalnev et al. (2024), such as SPARE Zhao et al. (2025) and SAE-SSV He et al. (2025).
6 Conclusion and Open Questions
In this survey, we offer a novel perspective on latent space geometry through the lens of compositional semantics, a direction we refer to as semantic representation learning. This direction enables a bridge between symbolic and distributional semantics, helping to mitigate the gap between them. By integrating structured semantic knowledge into learned representations, it holds the potential to address core challenges in machine learning—such as mechanistic interpretability, composition and generalisation. We review and compare three mainstream autoencoder architectures—VAE, VQVAE, and SAE—and examine the distinctive latent geometries they induce in relation to semantic structure and interpretability. Despite promising progress, several open questions remain:
Composition and Generalisation in NLI.
How can semantic representation learning improve compositional generalisation in high-level reasoning tasks such as Natural Language Inference (NLI)? Recent work by Fu and Frank (2024) explores how models can better generalise across compositional structures, but a more principled semantic grounding in the latent space remains an open challenge.
Compositional Mechanism Interpretability.
Whether autoregressive LLMs can exhibit human-interpretable compositional behaviours in their latent space—particularly in the context of Natural Language Inference (NLI)—remains an under-explored research area. While these models have demonstrated impressive performance on NLI benchmarks, their internal mechanisms for handling compositional reasoning are often opaque.
Diffusion-based LLMs.
Emerging diffusion-based LMs, such as Neuro-symbolic diffusion Nie et al. (2025); van Krieken et al. (2025b), offer new opportunities to combine the generative strengths of large-scale models with symbolic-level controllability. However, the design of semantically interpretable latent priors within these frameworks is still in its early stages. Incorporating the compositional semantic representation can potentially improve the sampling efficiency during the denoising process. Developing structured latent representations that align with formal semantics could enable more transparent reasoning and controllable generation in future generative architectures.
Appendix A Additional Examples
Figure 4: Generation evaluation for VAE or VQVAE latent space. More results can be found in Zhang et al. (2024b) for VAE and Zhang et al. (2024c) for VQVAE.
Figure 5: Formal semantic structure of natural language definition used in Carvalho et al. (2023).